OpenStudy (anonymous):

Brandon and Chris together can do a piece of work in 12 days. After Brandon has worked alone for 5 days, Chris finishes the work alone in 26 days. In what time can each alone do the work?

OpenStudy (anonymous):

we need some variables, two actually we can call bradon's rate \(x\) and chris's rate \(y\) then there combined rate is \(x+y\) and you know that they can complete one job in 12 days so we have one equation, namely \[12(x+y)=1\]

OpenStudy (anonymous):

the second thing we know is that \[5x+26y=1\] so we have two equations with two unknowns, \[12x+12y=1\]\[2x+26y=1\] and we can solve that system of equations for \(x\) and \(y\)

OpenStudy (anonymous):

sure did didn't i? glad you are here to keep me honest

OpenStudy (anonymous):

lets try \[12x+12y=1\] \[5x+26y=1\] maybe that will work better @hero do you have a better method? this is kind of clunky

OpenStudy (anonymous):

so i have to find solution to x and y now?

OpenStudy (anonymous):

ack, too bad, i hate these fractions at any rate solving gives \(x=\frac{1}{18}\) and \(y=\frac{1}{26}\) so brandon can do it in 18 hours and slow poke chris takes 36 hours

OpenStudy (anonymous):

another typo \(y=\frac{1}{36}\) i bet there is a snappier way

OpenStudy (anonymous):

@lynn17 yes you have to solve the two equations multiply the first one by 5, the second one by -12 and add, the \(x\) will drop out

OpenStudy (anonymous):

yeah other ways are just as messy, actually more so

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

how did you get from the fractions to the solution?

OpenStudy (anonymous):

from fractions to solutions if \(x=\frac{1}{18}\) then the rate is \(\frac{1}{18}\) i.e. you can do \(\frac{1}{18}\) of the job in one hour therefore it takes you 18 hours to do the job i.e. take the reciprocal